OpenStudy (anonymous):

Brooke bought a condominium for $342,700. She made a 12% down payment and financed the remaining amount using a 30-year fixed-rate mortgage at 6.7%. The monthly payment is $1,946. Brooke will pay for one discount point, a 0.75% origination fee, the brokerage fee, state documentary taxes on the deed and the mortgage, and the intangible tax. Discount points equal 1% of the mortgage amount. Documentary stamp tax on deed is $0.70 per $100 or portion thereof. Documentary stamp tax on mortgage is $0.35 per $100 or portion thereof. Mortgage broker fee is $225 plus 6% of the mortgage amount. Intangible tax is 0.2% of the mortgage amount. What is the total amount of the principal, interest, down payment, and fees described for Brooke’s condominium?

jigglypuff314 (jigglypuff314):

Hello @bsf219 and Welcome to OpenStudy! :) Thank you for asking a Qualified Help question! Please give me a few minutes to read through you question :)

jigglypuff314 (jigglypuff314):

First let us find out how much money is the mortgage :) That would be the original amount Minus 12% of that number so What is 12% of 342700 = ?

OpenStudy (anonymous):

41124!

jigglypuff314 (jigglypuff314):

good! then 342700 - 41124 = [mortgage amount] = ?

jigglypuff314 (jigglypuff314):

oh yeah, also, that number would find one of the numbers you are looking for :) the 41124 would be the "Down Payment" amount :D

OpenStudy (anonymous):

301,576 is the mortgage amount? And the 41124 is the down payment?

jigglypuff314 (jigglypuff314):

yes and yes :) the mortgage amount can also be called the "Principal" amount in this case because it will be the number that all the percents and interest rates are based on :)

OpenStudy (anonymous):

so adding them together is 342,700. So how do I figure out the interest and fees to add to that?

jigglypuff314 (jigglypuff314):

Interest = (Principal) * (the interest percent) * (number of years)

jigglypuff314 (jigglypuff314):

so for this problem, Interest = 301576 * 0.067 * 30 = ?

OpenStudy (anonymous):

606,167.76?

jigglypuff314 (jigglypuff314):

yes, that's what I got :)

OpenStudy (anonymous):

Okay so what about the fees? And do I add 606167.76 to the 342,700 to work towards my answer?

jigglypuff314 (jigglypuff314):

Nope, I don't believe you need to add your interest amount to anything for you answer :) for the fees, you would find them each separately then add them all together :) Discount Points = 1% of mortgage = 0.01 * 301576 = ? Origination Fee = 0.75% = 0.0075 * 301576 = ?

OpenStudy (anonymous):

3015.76 for discount points and 2261.82 for origination fee. But what about all of the taxes and intangible taxes?

jigglypuff314 (jigglypuff314):

those are right so far :) and I was getting to the others xD "Documentary stamp tax on deed is $0.70 per $100 or portion thereof. Documentary stamp tax on mortgage is $0.35 per $100 or portion thereof. Mortgage broker fee is $225 plus 6% of the mortgage amount. Intangible tax is 0.2% of the mortgage amount." although I am not sure how much the deed costs?

jigglypuff314 (jigglypuff314):

Documentary Stamp Tax on Mortgage = (0.35/100) * 301576 = ? Mortgage Broker Fee = 225 + (0.06 * 301576) = ? Intangible Tax = 0.002 * 301576 = ?

OpenStudy (anonymous):

1055.516 for doc stamp tax 18,319.56 for broker fee 603.152 for intangible tax. Add those all together and you get 19,978.228

OpenStudy (anonymous):

$370,354.71 $769,338.71 $772,752.00 $1,070,914.71 These are my answer choices

jigglypuff314 (jigglypuff314):

don't forget the two other fees from before :)

OpenStudy (anonymous):

adding in discount points/origination fees you get 367,955.808 which isnt one of my answer choices

jigglypuff314 (jigglypuff314):

I'll try to work it out ^_^

OpenStudy (anonymous):

Thank you!!

jigglypuff314 (jigglypuff314):

What numbers did you add to get 367,955.808 ? because I had missed out on the one part of adding the deed's fee amount since I did not know how much the deed was?

OpenStudy (anonymous):

I added the 342,700 (the downpayment & mortgage) + 19,978 (other taxes shown above) + 5,277.58 (points/orig fee)

OpenStudy (anonymous):

Its close to Option A which is 370,000 but i'm wondering if I missed something.

jigglypuff314 (jigglypuff314):

Yes, I am wondering if I read something wrong... another possibility would be 1946 monthly fee times 30 years times 12 months in a year = 700560 then Option A plus that number would get you Option D

OpenStudy (anonymous):

hmmm if you had to answer the question which one would you choose?

jigglypuff314 (jigglypuff314):

Option B because the 700560 is just the Interest plus the Mortgage that they want us to use which means adding the Down Payment and other Fees would get you a number very close to Option B

OpenStudy (anonymous):

Thank you so much!!

jigglypuff314 (jigglypuff314):

Glad I could help ^_^